# Diffusion Models for Time Series Forecasting: A Survey

### v1

@@ 5.1.2 Spatio-temporal Datasets
PEMS04 and PEMS08 [73]: Collected by CalTrans via the Performance Measurement System (PEMS) [74], these traffic datasets record 5-minute interval flow and speed data from loop detectors. PEMS04 contains readings from 307 sensors in California District 04 (Jan–Feb 2018), while PEMS08 includes data from 170 sensors in District 08 (July–Aug 2018).
Sea Surface Temperature (SST) [^10] https://www.psl.noaa.gov/data/gridded/data.noaa.oisst.v2.highres.html [75]: A daily-resolution global sea surface temperature dataset from NOAA OISSTv2.
Navier–Stokes Flow [76]: A synthetic fluid dynamics dataset simulating incompressible 2D flows on a 221×42 grid. Each instance includes randomly placed circular obstacles, with x/y velocity fields and pressure channels. It serves as a controlled benchmark for physical simulation and partial differential equation (PDE) forecasting.
Spring Mesh [76]: A synthetic physics dataset consisting of a 10×10 particle grid connected by springs. Each node has two positional and two momentum channels. The system evolves based on spring dynamics and provides a clean testbed for structure-aware sequence modeling.
@@ 5.2 Multimodal Datasets
Time-MMD (Multi-Domain Multimodal Dataset) [77]: The first large-scale multimodal time series dataset spanning nine diverse domains (e.g., agriculture, climate, economy, energy, health, traffic). Each sample includes aligned numerical time series and fine-grained textual context.
TTC (Time-Text Corpus) [78]: A benchmark comprising temporally aligned pairs of numerical time series and descriptive text in healthcare and climate science.
@@
TimeCAP (Time-Series Contextualization and Prediction) [79]: A framework where large language models serve as both contextualizers and predictors. The dataset includes real-world time series across domains (e.g., weather in New York, finance, healthcare test-positive rates), paired with generated summaries by GPT-4 and labeled event outcomes.
@@
Image-EEG [80]: A large-scale multimodal dataset that pairs time-resolved EEG signals with visual stimuli for human object recognition. It includes brain responses from 10 participants viewing over 22,000 images across 30 object categories.
Table 3: Performance comparison of five approaches on four datasets with a history length of 168 and averaged prediction length results.
Method Exchange Wind Electricity Weather
MSE MAE CRPS MSE MAE CRPS MSE MAE CRPS MSE MAE CRPS
RATD 0.013 0.073 0.339 0.784 0.579 0.673 0.151 0.246 0.373 0.281 0.293 0.301
TimeDiff 0.018 0.091 0.589 0.896 0.687 0.917 0.193 0.305 0.490 0.327 0.312 0.410
CSDI 0.077 0.194 0.397 1.066 0.741 0.941 0.379 0.579 0.480 0.356 0.374 0.354
mr-Diff 0.016 0.082 0.397 0.881 0.675 0.881 0.173 0.258 0.429 0.296 0.324 0.347
Table 4: Performance comparison of five approaches on six datasets with a history length of 96 and a prediction length of 192.
Method ILI Exchange Electricity Traffic ETTm2 Weather
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
TimeGrad 2.644 1.142 2.429 0.902 0.645 0.723 0.932 0.807 1.385 0.732 0.885 0.551
CSDI 2.538 1.208 1.662 0.748 0.553 0.795 0.921 0.678 1.291 0.576 0.842 0.523
SSSD 2.521 1.079 0.897 0.861 0.481 0.607 0.794 0.498 0.973 0.559 0.693 0.501
TimeDiff 2.458 1.085 0.475 0.429 0.730 0.690 1.465 0.851 0.284 0.342 0.277 0.331
TMDM 1.985 0.846 0.260 0.365 0.222 0.329 0.721 0.411 0.524 0.493 0.244 0.286
D3U 2.103 0.935 0.254 0.358 0.179 0.267 0.468 0.299 0.241 0.302 0.222 0.264
FALDA 1.666 0.821 0.165 0.296 0.163 0.248 0.412 0.251 0.246 0.301 0.215 0.255
Table 5: Performance comparison of six approaches on four datasets, where history and prediction window sizes are individually set: Exchange (90, 7), Weather (144, 144), Electricity (168, 24), and Solar (288, 144).
Method Exchange Weather Electricity Solar
NRMSE CRPS NRMSE CRPS NRMSE CRPS NRMSE CRPS
TimeGrad 0.066 0.010 0.691 0.527 0.043 0.024 0.973 0.416
SSSD 0.065 0.010 0.755 0.701 0.094 0.065 1.058 0.506
TimeDiff 0.074 0.012 0.702 0.635 0.045 0.036 0.882 0.423
TMDM 0.063 0.010 0.564 0.618 0.063 0.042 0.829 0.382
MG-TSD 0.075 0.009 0.696 0.596 0.032 0.019 0.756 0.375
StochDiff 0.052 0.008 0.491 0.521 0.048 0.029 0.812 0.391
@@ 6 Evaluation Metrics
The performance of a TSF system is evaluated from two complementary perspectives, namely, deterministic metrics and probabilistic metrics. Deterministic metrics compare a single point forecast to the observed value, while probabilistic metrics judge an entire predictive distribution. Tables 3 4 5 present the performance of several diffusion-based TSF methods on multiple benchmark datasets. In the following, we introduce and explain several commonly used evaluation metrics.
@@ 6.1 Deterministic Metrics
Deterministic metrics assess forecasting performance by computing point-wise errors between predicted and ground-truth values. Among them, Mean Squared Error (MSE) and its variants are widely used due to their sensitivity to large deviations and compatibility with gradient-based learning. Similarly, Mean Absolute Error (MAE) offers robust alternatives that are less affected by outliers and better suited for evaluating across different scales. These metrics provide intuitive and interpretable scores, making them a standard choice for deterministic time series forecasting. In the following, we provide detailed definitions and explanations of each metric.
Mean Squared Error (MSE)
is computed by
@@
Root Mean Squared Error (RMSE)
is computed by
@@
Normalized Root Mean Squared Error (NRMSE)
is computed by
@@
Mean Absolute Error (MAE)
is computed by
@@
where $\ell_{1}$–norm yields a piece-wise linear loss that is more robust to MSE.
Normalized Mean Absolute Error (NMAE)
is computed by
@@ 6.2 Probabilistic Metrics
Continuous Ranked Probability Score (CRPS)
is computed by
[…]
@@
Normalized Average CRPS (NACRPS)
To score an $H$-step, $d$-dimensional forecast, NACRPS aggregates the point-wise CRPS values and normalizes them by the total magnitude of the target window:
@@ 7 Discussions
In the aforementioned sections, we review the foundation diffusion models and diffusion-based TSF approaches, and then provide a systematic categorization and detailed description of these approaches, focusing on their conditioning sources and integration process. In this section, we summarize key characteristics of current work and discuss the main challenges and limitations that remain in diffusion-based TSF.
@@ 7.1 Overall Analysis of Existing Studies
Different from conditional image generation, where diffusion models primarily perform semantic feature transfer, diffusion-based TSF requires learning a temporal transformation from historical observations to future horizons. This fundamental difference introduces unique modeling challenges, prompting researchers to adapt the diffusion process to the temporal dynamics and data structure of time series. Many studies introduce domain-specific priors into the conditioning process. Preprocessing techniques such as trend-seasonality decomposition, multi-scale analysis, and frequency-domain transformations (e.g., Fourier decomposition) are widely adopted to disentangle structured components from the raw historical sequence. These extracted components, such as long-term trends, periodic cycles, and residuals, are incorporated into the diffusion model as structured conditioning inputs. This approach not only improves interpretability but also enhances the model’s ability to generate high-quality forecasts across different temporal scales and frequency bands. Another major direction incorporates multimodal conditioning. Compared with pure time series inputs, auxiliary modalities such as text or images offer additional semantic or contextual signals. For example, visual representations or textual descriptions aligned with temporal data guide the generation process toward more coherent or interpretable outputs. However, successful integration typically requires that these modalities are temporally synchronized and semantically relevant to the target series, which limits their general applicability in domains lacking such aligned data. In addition to enriching the conditioning source, several approaches focus on enhancing the way conditioning is integrated into the generative process. Rather than relying solely on static context vectors, recent approaches dynamically incorporate historical information through mechanisms such as adaptive noise scheduling, conditional priors embedded into the forward diffusion process, and reinterpreted diffusion trajectories that better reflect temporal causality. These designs seek to exploit the inherently sequential nature of both time series data and the diffusion process, enabling more controllable and temporally aligned generation. In summary, diffusion models hold great promise for TSF, and recent studies have made substantial progress in adapting them to temporal prediction tasks. These adaptations, including the use of temporal priors, multimodal conditioning, and customized integration mechanisms, reflect the increasing maturity of this research direction. Nevertheless, despite these advancements, diffusion-based TSF still faces a number of unresolved limitations and practical challenges.
7.2 Limitations of Current Research
The limitations mainly stem from three aspects. Firstly, TSF itself is inherently challenging. Secondly, diffusion models also come with certain natural limitations. Thirdly, there remains a lack of standardized evaluation protocols in the field.
Limitations from the Nature of TSF
TSF inherently presents several challenges that complicate the design of diffusion models. First, real-world time series often exhibit non-stationarity and abrupt regime shifts, making it difficult for models trained under stationary assumptions to generalize across varying temporal dynamics. Second, capturing long-term dependencies remains a fundamental obstacle, as the influence of distant past events may decay or interact in complex, nonlinear ways. Third, data sparsity and irregular sampling are common in practical settings such as healthcare or finance, where missing values and uneven time intervals severely hinder model performance. Lastly, multivariate time series often involve high-dimensional, heterogeneous variables with diverse temporal behaviors and interdependencies, posing significant challenges for stable joint modeling and representation learning.
Limitations of Current Diffusion Models
Current diffusion models for TSF also face several limitations. First, the iterative nature of the sampling process often leads to slow inference, requiring a large number of denoising steps to generate each forecast, which restricts their use in real-time applications. Second, most models rely solely on historical time series as the conditioning input, overlooking the potential benefits of incorporating richer modalities such as textual descriptions, spatial context, or external knowledge, which could enhance flexibility and generalization. Third, despite recent efforts [53, 51] to introduce interpretable components like trend or noise decomposition, the generation process remains largely opaque due to the use of high-dimensional latent spaces, limiting the interpretability of predictions. Finally, many existing models are tailored to fixed task settings and struggle to generalize across forecasting horizons, domains, or data distributions, indicating a lack of robustness and transferability in more realistic or zero-shot scenarios.
Lack of Standardized Evaluation Protocols
A notable limitation in current diffusion-based TSF research lies in the lack of standardized evaluation protocols. First, datasets are often directly sourced from public repositories and identified only by task-centric names such as “Traffic” or “Electricity”, without clear documentation of data preprocessing steps, variable selections, or temporal resolutions. This ambiguity hinders reproducibility and fair comparison. Second, there is substantial inconsistency in experimental setups across studies, particularly in the choice of historical window length and prediction horizon. These choices are often manually tuned per dataset or model, making cross-method comparisons difficult and potentially biased. Third, the field still lacks a unified benchmarking framework. Existing works adopt diverse metrics, evaluation pipelines, and test splits, which prevents a comprehensive and transparent comparison across approaches. Developing a standardized, modular evaluation benchmark remains an urgent need for future progress.

### v2

@@ 5.1.2 Spatio-temporal Datasets
  • PEMS04 and PEMS08141414https://pems.dot.ca.gov/[103]: Collected by CalTrans via the Performance Measurement System (PEMS) [104], these traffic datasets record 5-minute interval flow and speed data from loop detectors. PEMS04 contains readings from 307 sensors in California District 04 (Jan–Feb 2018), while PEMS08 includes data from 170 sensors in District 08 (July–Aug 2018).
  • Sea Surface Temperature (SST) [^15] https://www.psl.noaa.gov/data/gridded/data.noaa.oisst.v2.highres.html[105]: A daily-resolution global sea surface temperature dataset from NOAA OISSTv2. This high-resolution SST dataset provides daily, weekly, and monthly SST and sea ice data. As part of NOAA’s High-Resolution Blended Analysis, it features a 0.25 global grid covering 89.875S–89.875N and 0.125E–359.875E. The record spans September 1981 to August 2025 and includes a long-term monthly climatology (1991–2020) for reference.
  • Navier–Stokes Flow161616lhttps://archive.nyu.edu/handle/2451/63285[106]: A synthetic fluid dynamics dataset simulating incompressible 2D flows on a 221×42 grid. Each instance includes randomly placed circular obstacles, with x/y velocity fields and pressure channels. It serves as a controlled benchmark for physical simulation and partial differential equation (PDE) forecasting.
  • Spring Mesh171717lhttps://archive.nyu.edu/handle/2451/63285[106]: A synthetic physics dataset consisting of a 10×10 particle grid connected by springs. Each node has two positional and two momentum channels. The system evolves based on spring dynamics and provides a clean testbed for TSF approaches with structure-aware sequence modeling.
@@ 5.2 Multimodal Datasets
  • Time-MMD (Multi-Domain Multimodal Dataset) [^18] https://github.com/AdityaLab/Time-MMD/[19]: The first large-scale multimodal time series dataset spanning nine diverse domains (e.g., agriculture, climate, energy, and traffic). Each sample includes an aligned numerical time series and a fine-grained textual context.
@@
  • TimeCAP (Time-Series Contextualization and Prediction) [^19] https://github.com/geon0325/TimeCAP[107]: A framework where LLMs serve as both contextualizers and predictors. The dataset includes real-world time series across domains (e.g., weather in New York), paired with summaries generated by GPT-4 [108] and labeled event outcomes.
  • TTC (Time-Text Corpus) [^20] https://github.com/Rose-STL-Lab/Multimodal_Forecasting [109]: A benchmark comprising temporally aligned pairs of numerical time series and descriptive text in healthcare and climate science.
@@
  • Image-EEG [^21] https://osf.io/nb8wr[110]: A large-scale multimodal TSF dataset that pairs time-resolved electroencephalography (EEG) signals with visual stimuli for human object recognition. It includes brain responses from 10 participants viewing over 22,000 images across 30 object categories.
TABLE IV: MSE comparison across experimental settings with different history and prediction lengths. Underlined models denote original sources of reported metrics, while non-underlined entries represent concurrent baselines from those papers. Averages over length combinations are shown in braces {…}.
Dataset Exchange Weather Traffic Electricity ETTh1
History length {48,96,192,336}
Prediction length {96,168,336,720}
CCDM 0.4659 0.3414 0.8346 0.1864 0.4769
SSSD 0.7349 0.4742 0.9958 0.2532 1.0039
TimeDiff 0.4522 0.4316 0.9335 0.2048 0.4786
TMDM 0.8567 0.3275 0.9364 0.2047 0.5733
CSDI 0.7649 0.3065 1.4577 0.4012 1.0642
Dataset Exchange Weather Traffic Electricity ETTm2
History length {1000}
Prediction length {192}
TMDM 0.26 0.28 0.60 0.19 0.27
CSDI 1.67 0.86 0.94 0.56 1.28
SSSD 0.90 0.67 0.81 0.47 0.97
TimeDiff 0.48 0.36 0.68 0.27 0.41
Dataset Exchange ETTh1 ETTh2 ETTm1 ETTm2
History length {96}
Prediction length {96}
ARMD 0.093 0.445 0.311 0.337 0.181
MG-TSD 0.396 1.096 0.295 0.690 0.202
Dataset Exchange ETTh1 ETTh2 ETTm1 ETTm2
History length {96}
Prediction length 14 168 168 192 192
FDF 0.0204 0.4364 0.3393 0.3532 0.2303
CSDI 1.3023 1.1090 2.0700 1.0770 1.6051
TABLE V: MAE comparison across experimental settings with different history and prediction lengths. Underlined models denote original sources of reported metrics, while non-underlined entries represent concurrent baselines from those papers. Averages over length combinations are shown in braces {…}.
Dataset ILI Exchange Electricity Traffic ETTm2 Weather
History length {96}
Prediction length {192}
FALDA 0.821 0.296 0.248 0.251 0.301 0.255
CSDI 1.208 0.748 0.795 0.678 0.576 0.523
SSSD 1.079 0.861 0.607 0.498 0.559 0.501
TimeDiff 1.085 0.429 0.690 0.851 0.342 0.331
TMDM 0.846 0.365 0.329 0.411 0.493 0.286
D3U 0.935 0.358 0.267 0.299 0.302 0.264
Dataset ETTh1 ETTh2 ETTm1 ETTm2 ILI Solar
History length {168}
Prediction length 192 192 192 192 36 192
NsDiff 0.594 0.514 0.488 0.281 2.846 0.242
CSDI 0.949 1.226 1.002 1.723 4.515 0.763
TimeDiff 0.517 0.456 0.537 0.268 3.958 0.821
TMDM 0.696 0.512 0.494 0.315 3.636 0.250
Dataset Traffic Electricity Weather Exchange ETTh1 ETTm1
History length {96, 192, 336, 720, 1440}
Prediction length 168 168 672 14 168 192
mr-diff 0.197 0.332 0.032 0.094 0.196 0.149
CSDI 0.468 0.540 0.037 0.200 0.221 0.170
TimeDiff 0.207 0.341 0.035 0.102 0.202 0.154
SSSD 0.226 0.403 0.041 0.118 0.250 0.169
TABLE VI: CRPS comparison across experimental settings with varying history and prediction lengths. Underlined models denote the original source of reported metrics, while non-underlined entries represent concurrent baselines from those papers. Averages over length combinations are shown in braces {…}.
Dataset Exchange Wind Electricity Weather
History length {168}
Prediction length {96,192,336}
RATD 0.339 0.673 0.373 0.301
CSDI 0.397 0.941 0.480 0.354
TimeDiff 0.589 0.917 0.490 0.410
mr-diff 0.397 0.881 0.429 0.347
Dataset Solar Electricity Traffic taxi wiki
History length 168 168 168 48 90
Prediction length 24 24 24 24 30
CSDI 0.338 0.041 0.073 0.271 0.207
Dataset ETTh1 ETTh2 ETTm1 ETTm2 ILI
History length {168}
Prediction length 192 192 192 192 36
NsDiff 0.392 0.358 0.346 0.256 0.806
CSDI 0.492 0.647 0.524 0.817 1.244
TimeDiff 0.465 0.471 0.464 0.316 1.153
TMDM 0.452 0.383 0.375 0.289 0.967
Dataset Exchange Weather Traffic Electricity ETTh1
history length {48,96,192,336}
prediction length {96,168,336,720}
CCDM 0.3271 0.2563 0.3762 0.2111 0.3600
SSSD 0.4666 0.3118 0.4847 0.2630 0.5531
TimeDiff 0.4534 0.3526 0.5744 0.3169 0.4418
TMDM 0.5144 0.2791 0.5773 0.3163 0.4526
CSDI 0.5051 0.2344 0.6104 0.3069 0.5941
@@ 6 Evaluation Metrics
The performance of a TSF system is evaluated from two complementary perspectives, namely, deterministic metrics and probabilistic metrics. Deterministic metrics compare a single point forecast to the observed value, while probabilistic metrics judge an entire predictive distribution. Tables IV, V, and VI present the performance of several diffusion-based TSF approaches on multiple benchmark datasets. The following subsections illustrate widely used evaluation metrics.
@@ 6.1 Deterministic Metrics
Deterministic metrics assess forecasting performance by computing point-wise errors between predicted and ground-truth values. Among them, Mean Squared Error (MSE) and its variants are widely used due to their sensitivity to large deviations and compatibility with gradient-based learning. Similarly, Mean Absolute Error (MAE) offers robust alternatives that are less affected by outliers and better suited for evaluating across different scales. These metrics provide intuitive and interpretable scores, making them a standard choice for deterministic time series forecasting. In the following content, we provide detailed definitions and explanations of each metric.
Mean Squared Error (MSE) is computed by
@@
Root Mean Squared Error (RMSE) is computed by
@@
Normalized Root Mean Squared Error (NRMSE) is computed through the following process
@@
Mean Absolute Error (MAE) is computed by
@@
where $\ell_{1}$–norm yields a piece-wise linear loss that is more robust than MSE.
Normalized Mean Absolute Error (NMAE) is computed through the following process
@@ 6.2 Probabilistic Metrics
Continuous Ranked Probability Score (CRPS) is computed through the following process
$\operatorname{CRPS}(F,x)=\int_{-\infty}^{\infty}\left(F(z)-\mathbbm{1}\{x\leq z\}\right)^{2}\,dz$ (57)
@@
Normalized Average CRPS (NACRPS) To score an $H$-step, $d$-dimensional forecast, NACRPS aggregates the point-wise CRPS values and normalizes them by the total magnitude of the target window:
@@
Figure 5: The complex dependencies in time series. In the top panel, segments with sharp fluctuations are highlighted with a red background, where darker shades correspond to greater volatility and typically indicate more intricate temporal dynamics. In the bottom-left panel, four variables that are most correlated with the target variable (OT) are overlaid to demonstrate cross-variable coupling. The bottom-right panel presents the pairwise correlation matrix across variables, which reveals complex inter-variable dependencies.
Figure 6: Dynamic uncertainty in time series. In the top panel, the evolution of prediction intervals (PIs) and predictive variance highlights the time-varying nature of uncertainty. In the bottom panel, the PIs for four variables in a multivariate time series are shown, demonstrating how uncertainty evolves over time across different variables.
@@ 7 Discussions
In the aforementioned sections, we review the foundation diffusion models and diffusion-based TSF approaches, and then provide a systematic categorization and detailed description of these approaches, focusing on their conditioning sources and integration process. In this section, we summarize key characteristics of current work and discuss the limitations and challenges that remain in diffusion-based TSF.
@@ 7.1 Overall Analysis of Existing Studies
Key characteristics Diffusion-based TSF establishes a unified generative paradigm that addresses four core needs of temporal prediction. First, it models complex dependencies along the time axis and across variables through stepwise denoising with learnable score functions. This structure supports long range temporal effects and preserves cross time series interactions during sampling, consistent with the correlation patterns shown in Fig. 5. Second, it delivers probabilistic forecasts by drawing trajectories from the learned reverse dynamics, which yields full predictive distributions rather than only point values. The resulting intervals vary with regime changes and reflect dependence among variables, in agreement with the variability illustrated in Fig. 6. Third, the conditional formulation provides flexibility for injecting prior knowledge from the history and from auxiliary sources. Domain preprocessing separates structure from noise through trend and seasonality decomposition, multiscale analysis, and frequency domain transforms such as Fourier decomposition. The extracted components supply explicit conditions that guide central tendency and dispersion during denoising. Extensions broaden the source of conditions with textual context, spatial or relational graphs, and retrieval based references so that external cues complement the history when evidence in the past window is weak. Fourth, empirical studies on common benchmarks report competitive accuracy together with calibrated predictive intervals, and robustness often improves when conditions encode strong time series priors and when integration preserves temporal structure throughout the trajectory.
Two-layer perspective across approaches Existing approaches organize these gains around two coordinated layers. The first layer enriches the conditioning source so that the diffusion model receives disentangled temporal features and auxiliary evidence derived from the historical sequence and from complementary modalities. Representative choices include decomposition into trend, seasonality, and residual, multiscale summaries that capture coarse to fine patterns, and frequency domain components that describe periodic structure. Multimodal additions supply side information that clarifies context when the raw sequence is ambiguous and they expand the prior beyond what the history alone supports. The second layer refines the integration pathway so that conditions shape generation in a manner aligned with temporal dynamics and with consistent propagation of uncertainty. Some approaches inject conditions through the denoising network to steer the score toward patterns supported by priors, while other approaches embed conditions into the diffusion trajectory by adjusting schedules or latent covariances so that guidance acts within the forward and reverse dynamics. This two-layer view links observed advantages in dependency modeling, uncertainty estimation, conditioning flexibility, and benchmark performance to concrete choices on where priors enter and how they interact with the generative process, and it aligns the narrative with Fig. 5 and Fig. 6 that motivate the focus on correlations and dynamic uncertainty.
7.2 Limitations and Challenges
Model rigidity and generalization Many diffusion-based TSF approaches bind the forecasting specification to the architecture. The history length, the variable set, and the horizon are fixed by the input and output signature and by positional encodings, which restricts adaptation when any of these quantities changes, when channels appear or disappear, or when sampling becomes irregular. Structural priors optimized for a single configuration provide limited support for variable rate resampling and for missing segments, and they constrain the conditioning pathway to a narrow family of shapes. Retuning for each setting raises operational overhead and weakens reproducibility because hyperparameters and early stopping criteria drift across configurations. The absence of shape robust encoders and horizon aware heads also limits cross dataset transfer and out of domain generalization under distribution shift, since the learned score functions specialize to a restricted dependency range and to a fixed cross variable topology. The constraint further affects hierarchical forecasting and rolling updates where the specification must adjust on the fly, and it reduces reliability when new sensors appear, when some channels become inactive, or when covariates follow a different sampling regime.
Uncertainty estimation and interpretability Most approaches produce intervals or full predictive distributions, yet the mapping from interval width to causal factors remains unclear. Forecast require explanations of where risk concentrates and why a given step on the horizon receives wider bounds. Current pipelines rarely partition uncertainty into aleatoric and epistemic components or attribute variance to trend change, seasonal phase shift, exogenous shocks, or sensor noise. Calibration practice often centers on aggregate scores such as CRPS and underreports horizon wise reliability, tail behavior, and dependence across variables. Extreme events and regime transitions stress these weaknesses because heavy tailed errors and non stationary volatility receive little attention in the noise schedule or in the structure of the score network. Numerical validity therefore provides limited decision value when the intervals lack traceability to concrete signals in the history or in auxiliary inputs. In addition, interval sets sometimes violate monotonicity across confidence levels, quantiles cross on long horizons, and uncertainty fails to remain temporally coherent, which reduces trust in downstream risk controls.
Multimodal utilization and temporal asynchrony Recent approaches extend the conditioning source with text, spatial graphs, images, or external knowledge, yet integration frequently reduces to side feature concatenation to the denoiser. Multimodal information rarely shapes the forward diffusion prior or the reverse dynamics, which limits the influence of external context on generation. Streaming environments introduce asynchronous arrivals where the time series modality appears first and textual or knowledge signals arrive with latency. Without explicit alignment operators the model ignores delayed inputs or attaches them to the wrong forecast step, which injects noise that propagates across the trajectory and degrades both accuracy and calibration. Few studies explore asynchronous supervision during training, lag aware conditioning at inference, or memory mechanisms that revise beliefs once a delayed modality becomes available. The gap is larger when modalities are intermittently missing or when time stamps drift across data sources, since the alignment error accumulates across the sampling path.
Modality confidence and reliability weighting Heterogeneous sources present variable reliability over time. Textual reports may contain uncertainty markers, sensor nodes may drift, and retrieved references may mismatch context. Existing approaches assign static fusion weights or rely on attention without a calibrated notion of source confidence. The absence of confidence aware fusion affects the mean forecast and the spread of the distribution. Effective fusion requires reliability signals tied to the diffusion state or to the scheduler so that unreliable modalities receive reduced influence in high variance regimes and authoritative cues receive higher influence during regime transitions, which remains underexplored. Failure to manage reliability also hampers detection of modality dropouts and adversarial noise, and it prevents selective abstention when external evidence conflicts with the historical pattern.
Computational efficiency and deployability Iterative sampling introduces latency that grows with horizon and dimensionality. Long sequences demand more denoising steps for stability, and multivariate targets increase memory traffic and compute per step. High cost restricts real time deployment and narrows the scope of hyperparameter exploration during development. Accelerated solvers and distillation reduce step counts, yet aggressive compression risks calibration drift and distortion of cross variable dependence, which creates a trade off between speed and probabilistic fidelity. Systems factors add friction, including suboptimal batching for variable length windows and bandwidth bottlenecks during multistep sampling. Training cost also remains high because score models require long schedules, large batches for stable variance estimation, and frequent checkpointing, which limits iteration speed in applied settings.
Evaluation breadth and practical fidelity Benchmark coverage often concentrates on a limited set of datasets and fixed horizons, which hides sensitivity to configuration changes and to non stationary regimes. Many studies prioritize point metrics and a small set of probabilistic scores while omitting diagnostics that reveal horizon wise calibration, tail risk, conditional coverage across contexts, and stability of prediction sets under perturbations. Protocols rarely stress asynchronous multimodal input, delayed feedback, or realistic missingness patterns. Limited evaluation breadth obscures failure modes and slows progress on the limitations described above. Reporting sometimes omits computational budgets, solver step counts, and memory profiles, which complicates fair comparison and makes deployment trade offs less transparent.
Summary of open problems The field faces a linked set of obstacles. Structural adaptability remains limited by fixed input output shapes and narrow configuration choices, which reduces reliability under specification drift. Uncertainty estimation lacks interpretability that traces risk to trend, seasonality, exogenous events, or sensor reliability, and temporal coherence of intervals remains fragile. Multimodal fusion stays shallow and overlooks asynchrony and confidence, which reduces the value of external context in streaming environments. Computational cost escalates with sequence length and dimensionality, and naive acceleration threatens calibration and cross variable dependence. Evaluation practice trails practical needs and does not reflect deployment constraints such as latency budgets and intermittent modalities. Addressing these issues requires architectures that adjust structure without full retraining, uncertainty pathways that attribute risk to identified sources, fusion mechanisms that handle delays and reliability, sampling strategies that reduce steps without eroding probabilistic quality, and benchmarks that stress the scenarios encountered in real systems.